At what angle above the horizon is the sun if light from it is completely polarized upon reflection from water
1
Expert's answer
2020-11-24T05:52:54-0500


In order to be completely polarized, the incident angle (the angle with perpendicular to the surface) should be equal to Brewster's angle:


"\theta_B = \arctan\left( \dfrac{n_2}{n_1}\right)"

where "n_2 = 1.33" is the refractive index of water and "n_1 = 1" is the refractive index of air.

Thus, obtain:


"\theta_B = \arctan\left( \dfrac{1.33}{1}\right) \approx 53.1\degree"

The angle abov the horizon is then:


"\alpha = 90\degree - \theta_B = 36.9\degree"

Answer. "36.9\degree".